Shutdown activities

Infrastructure

TASK DATE COMMENTS STATUS
SG8 cooling regulation for RICH1 09/2009 (TBC) First proposal received from CV  
Cooling Towers maintenance weeks 18 to 20 Detailed schedule and cooling power availibility Done
Cooling DSS Flowswitches Change during Cooling Shutdown FS(low flow) for Muon A and C sides, Balcony, D3a -- Diff pressure sensors for UXA~C Mixed and Chilled? water Done
Detector cooling repair 21/11/08 Repair and clean flow meters, fix pneumatic leak, change PLC UPS, Change Backup chiller pump Done
Detector cooling annual maintenance week 18 to 20   Done
Gas systems maintenance 25/11 to 28/02 Up-to-date schedule Done
Access System maintenance 2-6/03 Includes Mobile plugs. Access problem limited to UXB on 09/3 Schedule and Impact Done
18 kV transformers conformity work Week 19 (7-8/05) Agreed with TS/EL Done
18 kV transformers triennial maintenance Week 19 (7-8/05) Agreed with TS/EL Done
18kV transformers coupling switch   Offer from TS/EL. Rejected.  
Mixed/Chilled water sand filters Week 26 Washing water recycling work remaining Done
Mixed water Standalone unit Week 18 DAI issued Done
Standalone cooling plants (Online, SG) maintenance 15-16/12/2008 Offer received. Done
Ventillation - Shielding against air from the tunnel     Ongoing since week 18
EXD/EOD UPS ventillation   Blowers to install  
Gas supply: C4F10 pipe heating blankets   Need to be checked / refurbished  
Gas supply monitoring      
Lifting Equipments control 10/02 - 12/02 All lifting tools concerned See list in logbook Done
UPS UX85 maintenance 19/2/2009 Report available Done
UPS UX85 pieces replacment No date yet Anticipation of 2010 replacements (batteries and other components will reach end of life in 2010)  
Refurbishment of Machine Safe Power for TFC 06/2009 Including AUL Done

Safety Systems

TASK DATE COMMENTS STATUS
Intercom XCR Week 34 Safe Communication with FB CCC Done
SNIFFER IS37 9/02 to 13/02 System off during day, operational in the evening Done
Evacuation sirens test 1h, to be scheduled (1-5/12/08?) Cavern to be evacuated Canceled
Hi-Fog / Marioff 4/12/08 Date confirmed by CMS. Report EDMS 980670 Done
Hi-Fog / Marioff 15/09/09 2009-2010 Maintenance  
AUG layout refurbishment and installation Completed week 35   Done
AUG tests 07/05/2009 Morning underground afternoon surface. Pit closed Done
DSS Maintenance and tests 07-08/05/2009   Done
Biometry PZ85 11/12/2008 Will be activated when final set of bypass padlock and keys will come Done
Biometry by-pass PZ and UX 14/11/08 Reception 17/11/08 Padlocks must be changed (final closure plan) Done
MAD key in XCR Padlock must be changed 14/11/08 Reception 17/11/08 Padlocks must be changed (final closure plan) Done
Flammable gas IS37 reception 11/12/2008 - 22/01/09 Reception 11/12/08 and 22/01/09 waiting for report from SC Done
CO2 (rack) system Test 09/01 and 23/01 Full test wit SC waiting for report from SC Done
ODH maintenance and IS37 yearly tests 09/02 to 13/02 Sirens and Flash light done during Xmas Done
Red Phones IS37   Request done to TS/ASE ongoing
Smoke UX85 maintenance and IS37 yearly test 16/2 to 20/2   Done
Smoke Racks UX85 maintenance 02/2 to 06/2   Done
Sump pump IS37 Week 18-20 Need confirmation from EN/CV Done
Evacuation IS37 test and maintenance + ODH link Xmas period 30/12/2008 Done
Evac button in XCR   Following task force recommendation Done
OT Sprinkler system Week 33 and Week 41 First part completed  
SX8 labs evacuation Tower 2009 Start week 9 Done

-- EricThomas - 21 Nov 2008

Edit | Attach | Watch | Print version | History: r38 < r37 < r36 < r35 < r34 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r38 - 2009-09-14 - EricThomas
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback